### v2

@@ 5 The impact of numerical chaos
In general, numerical integrations are constrained by the inherent limitations of floating-point arithmetic precision, the temporal discretization of the solution, and the truncation error of the integration algorithm (see Portegies Zwart & Boekholt, 2018 for a discussion). The simulations presented thus far also suffer from these limitations (even though in tsunami the truncation errors are compensated by employing Richardson extrapolation, Press et al., 2007). Small integration errors accumulate over time and result in a deviation of our approximate solution from the mathematical one (i.e., the true mathematical solution that is guaranteed to exist by virtue of the Cauchy–Lipschitz theorem), indicating a lack of numerical convergence. Previous studies have highlighted that as long as numerical errors are below a certain threshold (impossible to determine a priori), an ensemble of approximate solutions can be statistically indistinguishable, even when individual solutions may not all be numerically converged (Boekholt & Portegies Zwart, 2015).
@@
Here, we have the opportunity to quantify how numerical errors manifest in the initial phase space, in addition to understanding their impact on the statistical outcome of the 3BP. We anticipate that the effect of numerical errors will be more pronounced in the set with zoom level 14, the last set of Figure 5, where the separation of neighbouring initial conditions in phase space is the smallest.
To investigate this, we rerun the simulations using the brutus code (Boekholt & Portegies Zwart, 2015), which employs arbitrary-precision arithmetic. We set the error tolerance parameter to 10, which is more stringent by 17 orders of magnitude compared to tsunami, and a 256 bits floating point arithmetic, in place of the commonly used double-precision (64 bits).
In Figure 9, we compare the phase space of outcome states between the two integrators. Owing to its enhanced accuracy and precision, brutus resolves regular regions with higher resolution. The impact of numerical errors is evident in artificially mixing the phase space, diminishing the coherence and definition of regular regions. Fine regular stripes resolved in the bottom panel lose coherence at lower accuracy and blend together with neighboring stripes. We term this artificial mixing as numerical chaos.
Since the term “numerical chaos” has only been introduced once before in the literature (Ablowitz et al., 1993), it is essential to provide a clear definition to prevent confusion. Numerical chaos arises when numerical errors are amplified within a system characterized by both regular and chaotic regions in its phase space. For numerical chaos to occur, chaotic regions must exist, which diffuse throughout the phase space and infiltrate otherwise regular regions because of numerical errors. Put simply, numerical errors cannot induce chaos in a system that is entirely regular.
@@ 6 Perturbations below the Planck scale
With compelling evidence demonstrating that chaos and regularity persist at any scale, including scales below the Planck length, any microscopic stochastic effect below the Planck length has the potential to propagate macroscopically (see also Boekholt et al., 2020). In this context, the regular behaviour of the 3BP provides a means to connect micro- and macroscopic physics.
@@ 7 Astrophysical implications
Our findings raise concerns regarding the application of statistical escape theories to astrophysical scenarios. While all trajectories we calculated numerically are physically plausible and can occur in nature, statistical escape theories only predict those trajectories that exhibit sufficient chaotic behavior to mix the phase space. In our setup, this results in the exclusion of approximately 32% of the initial phase space. As demonstrated in Section 3, this exclusion biases the distribution of outcome parameters (such as binary semimajor axis and eccentricity, and single escape velocity), as well as the escape probability of the single, based on its mass.
@@
In Figure 11, we illustrate how these biases impact the estimation of the gravitational wave coalescence time for the final binaries in our simulations, assuming they are black holes (Peters, 1964). Despite constituting only about a third of the total, binaries formed from regular interactions exhibit significantly shorter coalescence times compared to those formed from chaotic interactions, because of their higher eccentricity. Moreover, binaries resulting from regular interactions with short () coalescence times are roughly twice as numerous as those from chaotic interactions. Our analysis therefore suggests that statistical escape theories overestimate the coalescence time, and thus underestimate the merger efficiency of binaries formed from three-body interactions. Another likely consequence of the bias in eccentricity is the underestimation of the fraction of eccentric mergers that may be detectable with gravitational wave interferometers (Abbott et al., 2019; Romero-Shaw et al., 2021; Gayathri et al., 2022; The LIGO Scientific Collaboration et al., 2023).
@@
We emphasize that the fraction of regular trajectories, as well as their outcome properties, should vary depending on the specific initial setup of the 3BP. Nevertheless, we anticipate the coexistence of regular and chaotic regions to be a natural feature of the 3BP, as supported by findings in other studies with different setups (e.g. Manwadkar et al., 2021; Parischewsky et al., 2023; Manwadkar et al., 2023). For example, Manwadkar et al. (2023) investigated the ’chaotic absorptivity’ (the probability for an interaction to enter chaotic motion) for hyperbolic binary-single encounters with varying initial binary angular momentum and energy. Their figure 5 illustrates the persistent separation between chaotic and regular motion in the energy-angular momentum space, with the fraction of chaotic phase space decreasing for larger initial binary separations. Thus, it is important to evaluate the biases introduced by the use of statistical escape theories to model astrophysical scenarios on a case-by-case basis. A comprehensive assessment of these biases is beyond the scope of this paper.
@@ 8 Conclusions
Our numerical investigation highlights the mixed nature of chaos in the gravitational 3BP. At all scales, chaos is intermixed with regular regions where the triple breaks up very quickly, exhibiting regular behaviour, identified as a smooth mapping between initial conditions and outcome properties (Figure 3). These regular interactions comprise about 32% of the total phase space.
@@
Eventually, chaotic behaviour reaches the quantum realm, where a difference in positions below the Planck length results in macroscopic different outcomes (Figure 10). The mixing of regular and chaotic regions persists below this scale, coupling macro and microscopic physics.
@@
One potential critique of our work could be that our results are only applicable to our specific three-body setup, which admittedly represents a small portion of the vast parameter space of the 3BP. To address this concern, we conducted additional numerical experiments involving hyperbolic binary-single scatterings with an initial velocity at infinity and impact parameter. Remarkably, we observed the same regular patterns as depicted in Figure 3, indicating the robustness of our findings. Moreover, our conclusions are supported by the results of Manwadkar et al. (2023), who found that chaotic and regular motion occupy distinct regions in the energy-angular momentum space of hyperbolic binary-single encounters, even when statistically averaged over the binary phase.
@@
Our findings underscore that regular regions in the phase space pose challenges to the effectiveness of statistical escape theories as predictive tools for three-body interactions. Consequently, caution is advised when applying such tools to astrophysical problems (e.g. Kritos et al., 2022; Mapelli et al., 2022; Grondin et al., 2023; Ginat & Perets, 2023). Specifically, the eccentricity distribution of final binaries from regular trajectories diverges significantly from the expectations of statistical escape theories, which introduces a bias into gravitational wave population synthesis models reliant on such theories.
@@ Acknowledgements.
AAT acknowledges support from JSPS KAKENHI Grant Number 21K13914 and from the European Union’s Horizon 2020 and Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreements no. 847523 and 101103134. All the simulations were performed on the awamori computer cluster at The University of Tokyo. TCNB is supported by an appointment to the NASA Postdoctoral Program at the NASA Ames Research Center, administered by Oak Ridge Associated Universities under contract with NASA. NWCL gratefully acknowledges the generous support of a Fondecyt General grant 1230082, as well as support from Núcleo Milenio NCN2023_002 (TITANs) and funding via the BASAL Centro de Excelencia en Astrofisica y Tecnologias Afines (CATA) grant PFB-06/2007. NWCL also thanks support from ANID BASAL project ACE210002 and ANID BASAL projects ACE210002 and FB210003. AAT, SPZ and TCNB also thank Anna Lisa Varri for organizing the “Chaotic rendezvous” meeting at the University of Edinburgh, where part of this work was completed. AAT thanks Mario Pasquato for being an accessible source of statistical knowledge at any time of the day.References

### v3

@@ 5 The impact of numerical chaos
In general, numerical integrations are constrained by the inherent limitations of floating-point arithmetic precision, the temporal discretization of the solution, and the truncation error of the integration algorithm (see Portegies Zwart & Boekholt, 2018 for a discussion). The simulations presented thus far also suffer from these limitations (even though in tsunami the truncation errors are compensated by employing Richardson extrapolation, Press et al., 2007). Small integration errors accumulate over time and result in a deviation of our approximate solution from the mathematical one (i.e. the true mathematical solution that is guaranteed to exist by virtue of the Cauchy–Lipschitz theorem), indicating a lack of numerical convergence. Previous studies have highlighted that as long as numerical errors are below a certain threshold (impossible to determine a priori), an ensemble of approximate solutions can be statistically indistinguishable, even when individual solutions may not all be numerically converged (Boekholt & Portegies Zwart, 2015).
@@
Here, we have the opportunity to quantify how numerical errors manifest in the initial phase space, in addition to understanding their impact on the statistical outcome of the 3BP. We anticipate that the effect of numerical errors will be more pronounced in the set with zoom level 14, the last set of Figure 5, where the separation of neighbouring initial conditions in the phase space is the smallest.
To investigate this, we rerun the simulations using the brutus code (Boekholt & Portegies Zwart, 2015), which employs arbitrary-precision arithmetic. We set the error tolerance parameter to 10-30, which is more stringent by 17 orders of magnitude compared to tsunami, and a 256 bits floating point arithmetic, in place of the commonly used double-precision (64 bits).
In Figure 9, we compare the phase space of outcome states between the two integrators. Owing to its enhanced accuracy and precision, brutus resolves regular regions with higher resolution. The impact of numerical errors is evident in artificially mixing the phase space, diminishing the coherence and definition of regular regions. Fine regular stripes resolved in the bottom panel lose coherence at lower accuracy and blend together with neighboring stripes. We term this artificial mixing as ‘numerical chaos’.
Because the term numerical chaos has only been introduced once before in the literature (Ablowitz et al., 1993), it is essential to provide a clear definition to prevent confusion. Numerical chaos arises when numerical errors are amplified within a system characterized by both regular and chaotic regions in its phase space. For numerical chaos to occur, chaotic regions must exist, which diffuse throughout the phase space and infiltrate otherwise regular regions because of numerical errors. Put simply, numerical errors cannot induce chaos in a system that is entirely regular.
@@ 6 Perturbations below the Planck scale
With compelling evidence demonstrating that chaos and regularity persist at any scale, including scales below the Planck length, any microscopic stochastic effect below the Planck length has the potential to propagate macroscopically (see also Boekholt et al., 2020). While chaos washes away any microscopic effect by hiding the correlation between initial conditions and final outcomes, regular behaviour offers a means to connect micro- and macroscopic physics. Specifically, any potential microscopic perturbation to particle positions will not produce a different macroscopic outcome if the initial position resides within a regular region, as opposed to residing within a chaotic region.
@@ 7 Astrophysical implications
Our findings raise concerns regarding the application of statistical escape theories to astrophysical scenarios. While all trajectories we calculated numerically are physically plausible and can occur in nature, statistical escape theories only predict those trajectories that exhibit sufficient chaotic behaviour to mix the phase space. In our setup, this results in the exclusion of approximately 37% of the initial phase space. As demonstrated in Section 3, this exclusion biases the distribution of outcome parameters (such as binary semimajor axis and eccentricity, and single escape velocity), as well as the escape probability of the single, based on its mass.
@@
In Figure 11, we illustrate how these biases impact the estimation of the gravitational wave coalescence time for the final binaries in our simulations, assuming they are black holes (Peters, 1964). Despite constituting only about a third of the total, binaries formed from regular interactions exhibit significantly shorter coalescence times compared to those formed from chaotic interactions, because of their higher eccentricity. Moreover, binaries resulting from regular interactions with short () coalescence times are roughly three times as numerous as those from chaotic interactions. Our analysis therefore suggests that statistical escape theories overestimate the coalescence time, and thus underestimate the merger efficiency of binaries formed from three-body interactions. Another likely consequence of the bias in eccentricity is the underestimation of the fraction of eccentric mergers that may be detectable with gravitational wave interferometers (Abbott et al., 2019; Romero-Shaw et al., 2021; Gayathri et al., 2022; The LIGO Scientific Collaboration et al., 2023).
@@
We emphasize that the fraction of regular trajectories, as well as their outcome properties, should vary depending on the specific initial setup of the 3BP (see Appendix A). Nevertheless, we anticipate the coexistence of regular and chaotic regions to be a natural feature of the 3BP, as supported by findings in other studies with different setups (e.g. Manwadkar et al., 2021; Parischewsky et al., 2023; Manwadkar et al., 2023). For example, Manwadkar et al. (2023) investigated the ‘chaotic absorptivity’ (the probability for an interaction to enter chaotic motion) for hyperbolic binary-single encounters with varying initial binary angular momentum and energy. Their figure 5 illustrates the persistent separation between chaotic and regular motion in the energy-angular momentum space, with the fraction of chaotic phase space decreasing for larger initial binary separations. Thus, it is important to evaluate the biases introduced by the use of statistical escape theories to model astrophysical scenarios on a case-by-case basis. A comprehensive assessment of these biases is beyond the scope of this paper.
Another caveat is that our simulations are fully Newtonian and do not account for collisions between finite-sized particles. The inclusion of post-Newtonian (PN) corrections would introduce the possibility of mergers (or collisions), driven by radiation reaction terms (2.5 and 3.5 PN order corrections) and apsidal precession effects (1, 2, and 3 PN order corrections). However, the presence of dissipation terms is unlikely to significantly affect the mixing of chaos and regular regions. This is because the radiation reaction terms depend steeply on the distance between two particles, and are efficient only during the closest approaches. Previous studies have shown that radiation terms create narrow regular regions in the phase space, corresponding to collision outcomes, while leaving the rest of the phase space largely untouched (Samsing & Ilan, 2019; Parischewsky et al., 2023). On the other hand, precession terms have been found to reduce chaos in -body systems, but these effects are only significant in the strong gravity regime (, Portegies Zwart et al. 2022).
In summary, the inclusion of PN corrections would likely aggravate the bias of statistical escape theories by increasing the regularity of the system. We defer a complete numerical verification of this statement to future work, as we have disabled PN corrections for the present study to enable a direct comparison of our simulations with statistical escape theories, which assume Newtonian gravity.
@@ 8 Conclusions
Our numerical investigation highlights the mixed nature of chaos in the gravitational 3BP. At all scales, chaos is intermixed with regular regions where the triple breaks up very quickly, exhibiting regular behaviour, identified as a smooth mapping between initial conditions and outcome properties (Figure 3). These regular interactions comprise about 37% of the total phase space.
@@
Eventually, the chaotic behaviour reaches the quantum realm, where a difference in positions below the Planck length results in macroscopic different outcomes (Figure 10). The mixing of regular and chaotic regions persists below this scale, coupling macro and microscopic physics.
@@
One potential critique of our work could be that our results are only applicable to our specific three-body setup, which admittedly represents a small portion of the vast parameter space of the 3BP. To address this concern, we conducted additional numerical experiments involving hyperbolic binary-single scatterings with an initial velocity at infinity and impact parameter (see Appendix A). Remarkably, we observe regular patterns similar to those in Figure 3, and estimate the percentage of regular interactions to range from 28% to 84%, depending on the initial setup (Figure 12). These complementary results indicate the robustness of our findings. Moreover, our conclusions are supported by the results of Manwadkar et al. (2023), who found that chaotic and regular motion occupy distinct regions in the energy-angular momentum space of hyperbolic binary-single encounters, even when statistically averaged over the binary phase.
@@
Our findings underscore that regular regions in the phase space pose challenges to the effectiveness of statistical escape theories as predictive tools for three-body interactions. Consequently, caution is advised when applying such tools to astrophysical problems (e.g. Kritos et al., 2022; Mapelli et al., 2022; Grondin et al., 2023; Ginat & Perets, 2023). Specifically, the eccentricity distribution of final binaries from regular trajectories diverges significantly from the expectations of statistical escape theories (Figure 11), which introduces a bias into gravitational wave population synthesis models reliant on such theories.
@@ Acknowledgements.
The authors thank the editor Benoît Noyelles and the anonymous referee for suggestions that improved the paper. AAT acknowledges support from JSPS KAKENHI Grant Number 21K13914 and from the European Union’s Horizon 2020 and Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreements no. 847523 and 101103134. All the simulations were performed on the awamori computer cluster at The University of Tokyo. TCNB is supported by an appointment to the NASA Postdoctoral Program at the NASA Ames Research Center, administered by Oak Ridge Associated Universities under contract with NASA. NWCL gratefully acknowledges the generous support of a Fondecyt General grant 1230082, as well as support from Núcleo Milenio NCN2023_002 (TITANs) and funding via the BASAL Centro de Excelencia en Astrofisica y Tecnologias Afines (CATA) grant PFB-06/2007. NWCL also thanks support from ANID BASAL project ACE210002 and ANID BASAL projects ACE210002 and FB210003. AAT, SPZ and TCNB also thank Anna Lisa Varri for organizing the “Chaotic rendezvous” meeting at the University of Edinburgh, where part of this work was completed. AAT would like to express gratitude to those who contributed with insightful discussions at different stages of this work, including Alessandro Ballone, Rosemary Mardling, Barak Kol, Mario Pasquato.References
@@
Appendix A Simulations with hyperbolic setup
We performed additional simulations to investigate whether the mixed chaotic-regular phase space observed in the 3BP is a characteristic feature of our specific initial setup or not. In these new nine sets of simulations, we initialized the binary and the single on hyperbolic trajectories with a velocity at infinity and impact parameter . For each set, we run realizations, varying the binary phase and inclination of the incoming single, but fixing velocity at infinity and impact parameter. Specifically, we choose a combination of , and , where is the critical velocity for which the total energy of the three-body system vanishes:
and , and . The masses (, , ) and semi-major axis () of the binary are kept the same as in our fiducial experiment. We run realizations for each set.
Figure 12 shows the initial phase-space maps, akin to Figure 3, for each of the nine sets. Despite the radically different initial setups, we observe similar patterns to those in our fiducial setup, where the centre-of-mass velocities of the binary and the single are zero.
We calculated the fraction of regular interactions as in Section 3, employing neighbours. These fractions are depicted in Figure 12. Overall, the fraction of regular interactions ranges from to , and for small and (top-left panel) it is comparable to our fiducial setup. Notably, for and , the phase space closely resembles the one in Figure 3. However, as and increase, the phase-space maps begin to morph, displaying stronger asymmetries and more complex shapes. The fraction of regular interactions increases with , most likely due to the increase in flybys. Curiously, the fraction of regular interactions decreases for increasing impact parameter at .
Appendix B Zoom-ins in other regions
Figure 13 displays the initial phase space for alternative zoom-ins at levels 5, 6, and 7. This zoom-in targeted a region at the border between regular and chaotic regions. In the last two zoom-ins, the regular region occupies half of the box, while the chaotic region shows no resolved regular substructures. The measured fractal dimensions indicate the absence of complexity of the regular regions, with values of , , and for the three zoom levels, respectively. In the last two zoom levels, , reflecting the lack of resolved regular substructures within the chaotic region.
We speculate that with arbitrarily high resolution and numerical accuracy, it would be possible to resolve the regular regions within the chaotic region. Fine regular stripes can be observed on the left side of zoom level 5 (also visible in the outer 4th zoom in Figure 5), which progressively become finer until disappearing into the chaotic region on the right. These stripes likely correspond to commensurabilities between the return time of the single and the period of the leftover binary. As explained in Section 4, the regular regions are bordered by extremely long excursions, which can easily accumulate numerical errors due to the extended integration time. These numerical errors can then dephase the coherence of regular regions in phase space (as discussed in Section 5), creating the appearance of a chaotic region where regular substructures should instead be present.
These zoom-ins were discarded due to the extensive computational time required to resolve them, since they targeted the border of a regular region. As a result, the 5th zoom-in in Figure 5 was off-centred. Nevertheless, they illustrate the sensitivity of the fractal dimension measurement to the specific region of the initial phase space.